OpenStudy (anonymous):

(6 (7/8))(22(2/5))÷(7 (1/8))=

OpenStudy (anonymous):

u the eqution at the bottom rewite

OpenStudy (dumbcow):

convert to improper fractions first \[\frac{55}{8}*\frac{112}{5}=14*11=154\] \[\frac{154}{1}\div \frac{57}{8}=\frac{154}{1}*\frac{8}{57}=\frac{1232}{57}=21\frac{35}{57}\]